OpenStudy (anonymous):

A public swimming pool is rectangular with perimeter of 150 meters. The length of the pool is 50 meters less than four times its width. Determine the dimensions,in meters, of the pool.

OpenStudy (anonymous):

Ok, I got 25*4=100-50=50 50+50+25+25=150

OpenStudy (anonymous):

150 = P 4W = L+50

OpenStudy (anonymous):

So: Length: 50 Meters Width: 25 Meters

OpenStudy (anonymous):

Sovle the following for L and w. {2 (L+w)=150,L=4 w-50}

OpenStudy (anonymous):

@OrangeMaster is correct.

OpenStudy (anonymous):

Ok thank you @OrangeMaster and everybody else